# Gaussian Head Avatar: Ultra High-fidelity Head Avatar via Dynamic Gaussians

### v1

@@ Abstract
Creating high-fidelity 3D head avatars has always been a research hotspot, but there remains a great challenge under lightweight sparse view setups. In this paper, we propose Gaussian Head Avatar represented by controllable 3D Gaussians for high-fidelity head avatar modeling. We optimize the neutral 3D Gaussians and a fully learned MLP-based deformation field to capture complex expressions. The two parts benefit each other, thereby our method can model fine-grained dynamic details while ensuring expression accuracy. Furthermore, we devise a well-designed geometry-guided initialization strategy based on implicit SDF and Deep Marching Tetrahedra for the stability and convergence of the training procedure. Experiments show our approach outperforms other state-of-the-art sparse-view methods, achieving ultra high-fidelity rendering quality at 2K resolution even under exaggerated expressions. Project page: https://yuelangx.github.io/gaussianheadavatar/.
@@ 1 Introduction
To overcome this bottleneck and further improve the avatar quality, we introduce 3D Gaussian splatting [26] for 3D head avatar modeling. This is an explicit discrete representation that can be well adapted to the rasterization-based rendering pipeline. It has been verified that the 3D Gaussian representation is capable of render complex scenes with low computational consumption. Compared to NeRF, the reconstruction quality of static and dynamic scenes [57, 40, 65] is much better while rendering time cost has been significantly reduced. Motivated by this progress, we propose Gaussian Head Avatar, a novel representation that utilizes 3D Gaussian splatting for ultra high-fidelity head avatar modeling. Although recent 4D Gaussian works [57, 65, 40, 64] have been proposed to reconstruct dynamic scenes, all of them cannot be animated. For modeling the animatable head avatar, it is crucial but still unexplored how to effectively control the deformation of 3D Gaussians and model the dynamic appearances through expression coefficients.
@@ 2 Related Works
3D Head Avatar Reconstruction. Due to the wide application value in the film and digital huamn industry, 3D head avatar reconstruction from multi-view images has always been a research hotspot. Traditional works [32, 4, 19, 8] reconstruct the scan geometry through multi-view stereo and then register a face mesh template to it. However, such methods usually require heavy computation. With the utilization of deep neural networks, current methods [34, 7, 59, 63] achieve very fast reconstruction, producing even more accurate geometry. Lombardi et al. [37], Bi et al. [5] and Ma et al. [41] represent the full head mesh through a deep neural network and train it with multi-view videos as supervision. However, due to the errors in geometric estimation, mesh-based head avatars typically suffer from texture blur. Therefore, some recent methods [55, 38] utilize NeRF representation [43] to synthesize novel view images without geometry reconstruction, or build NeRF on the head mesh template [39]. Furthermore, the NeRF-based methods are extended to sparse view reconstruction tasks [69, 42, 47, 29] and achieve impressive performance.
@@
3D head avatars reconstruction from monocular videos is also a popular yet challenging research topic. Early methods [12, 13, 25, 24, 15, 45] optimize a morphable mesh to fit the training video. Recent methods [20, 27] leverage neural networks to learn non-rigid deformation upon 3DMM face templates [33, 18], thus can recover more dynamic details. Such methods are not flexible enough to handle complex topologies. Therefore, the latest methods explore to construct head avatar models based on implicit SDF [70], point clouds [71] or NeRF [21, 36, 16, 3, 2, 17, 61, 73, 62].
Point-based Rendering. Point elements as a discrete and unstructured representation can fit geometry with arbitrary topology [67] efficiently. Recent methods [56, 31, 30] open up a differentiable rasterization pipeline, such that the point-based representation is widely used in multi-view reconstruction tasks. Aliev et al. [1] and Ruckert et al. [48] propose to first render the feature map, which is transferred to the images through a convolutional renderer. Xu et al.[60] use neural point cloud associated with neural features to model a NeRF. Recently, 3D Gaussian splatting [26] shows its superior performance, beating NeRF in both novel view synthesis quality and rendering speed. Some approaches [57, 65, 40, 64] extend Gaussian representation to dynamic scene reconstruction. However, these methods can not be migrated to the head avatar reconstruction tasks.
@@ 3 Overview
The pipeline of the reconstruction of Gaussian Head Avatar is illustrated in Fig. 2. Before the beginning of the pipeline, we remove the background [35] of each image and jointly estimate the 3DMM model [18], 3D facial landmarks and the expression coefficients for each frame. In the initialization stage (Sec. 4.3), we reconstruct an SDF-based neutral geometry, and optimize a deformation MLP and a color MLP from the training data as the guidance model. Next, we extract the neutral mesh through DMTet to initialize the neutral Gaussians while the deformation and color MLPs are also inherited from the initialization stage. In the training stage of Gaussian Head Avatar (Sec. 4.2), we deform the neutral Gaussians to the target expression through the dynamic generator given the driving expression coefficients as the condition. Finally, given a camera view, the expressive Gaussians are rendered to a feature map, which is then fed into the convolutional super resolution network to generate high-resolution avatar images. The whole model is optimized under the supervision of multi-view RGB videos.
@@ 4.1 Avatar Representation
$\lambda_{exp}(\cdot)$ and $\lambda_{pose}(\cdot)$ represent the extent to which the points are affected by the expression or the head pose respectively. Here, we assume that the Gaussian points closer to 3D landmarks are more affected by the expression coefficients and less affected by the head pose, while the opposite is true for the Gaussian points far away. Specifically, The 3D landmarks $\boldsymbol{P}_{0}$ of the canonical model are first estimated through the 3DMM model in the data preprocessing and then optimized in the initialization stage 4.3. Then for each Gaussian point, we calculate the above weight $\lambda_{exp}(\cdot)$ and $\lambda_{pose}(\cdot)$ as follows:
@@ 4.2 Training
During training, we jointly optimize all the learnable paramters mentioned above in bold. For the loss function, we only use the foreground RGB images $I_{gt}$ as supervision to construct an L1 loss and a VGG perceptual loss [68] between the generated images $I_{hr}$ and the ground truth $I_{gt}$. Besides, we encourage the first three channels of the 32-channel feature image $I_{C}$ to be RGB channels, which is ensured by a L1 loss term. The total loss is:
@@
with $I_{lr}$ denoting the first three channels of the 32-channel image $I_{C}$. We set the weights $\lambda_{vgg}=0.1$ and $\lambda_{lr}=0.1$.
@@ 4.3 Geometry-guided Initialization
Due to the fact that the Gaussians representation is unordered and unstructured, it is difficult for the gradient to continueously spread to the nearby points in space after it is propagated back to one point. Consequently, randomly initializing the neutral Gaussians usually leads to failure to converge and initializing with a FLAME model fails to model the long hairstyle and the shoulders. To overcome this problem, we propose to utilize the implicit signed distance field (SDF) representation and Deep Marching Tetrahedra (DMTet) to first reconstruct a neutral mesh for initializing the Gaussian positions. We also coarsely optimize the color MLPs and the deformation MLPs in Sec.4.1 as well.
Representation and Rendering. Specifically, we first construct a MLP $\boldsymbol{f}_{sdf}$ to represent a signed distance field. In addition, this network will also output the corresponding feature vector of each point, which is used for predicting the point color. It can be formulated as:
@@
Parameters Transfer. Finally, we use the roughly trained guidance model to initialize the Gaussian head model. Specifically, we extract the neutral mesh with vertices $X$ and per-vertex features $F$ through DMTet, and directly assign their values to the neutral positions $\boldsymbol{X}_{0}=X$ and the per-vertex feature vectors $\boldsymbol{F}_{0}=F$ of the neutral Gaussians respectively. For the other neutral attributes, we adopt the original initialization strategy in Gaussian Splatting [26]. Then, we retain all the four optimized MLPs: $\{\boldsymbol{f}_{col}^{exp},\boldsymbol{f}_{col}^{pose},\boldsymbol{f}_{def}^{exp},\boldsymbol{f}_{def}^{pose}\}$ for the next stage, while the parameters of the two attribute MLPs: $\{\boldsymbol{f}_{att}^{exp},\boldsymbol{f}_{att}^{pose}\}$ and the super resolution network $\boldsymbol{\Psi}$ are randomly initialized.
@@ 5.1 Implementation Details
During the geometry-guided initialization stage, we use an Adam [28] optimizer, and set the learning rate to $1\times 10^{-3}$ for all the networks and $1\times 10^{-4}$ for the neutral 3D landmarks $\boldsymbol{P}_{0}$. Then We train the model for 10000 iterations with a batch size of 4.
During the Gaussian model training stage, we also use an Adam optimizer, and set the learning rate to $1\times 10^{-4}$ for the two color MLPs and two attribute MLPs, $1\times 10^{-5}$ for the two deformation MLPs, $1\times 10^{-5}$ for the neutral positions $\boldsymbol{X}_{0}$, $1\times 10^{-5}$ for the point-wise feature vectors $\boldsymbol{F}_{0}$, $1\times 10^{-4}$ for the neutral rotation $\boldsymbol{Q}_{0}$, $3\times 10^{-4}$ for the neutral scale $\boldsymbol{S}_{0}$, $1\times 10^{-3}$ for the neutral opacity $\boldsymbol{Q}_{0}$ and $1\times 10^{-4}$ for the super resolution network $\boldsymbol{\Psi}$. Finally, we train the Gaussian model for 600000 iterations with a batch size of 1 until fully convergence.
@@
Table 1: Quantitative evaluation results of NeRFBlendShape [17], NeRFace [16], HAvatar [69], our method without super resolution and our full method.
@@ 5.2 Results and Comparisons
In this section, we first compare our method with existing SOTA methods in qualitative experiments on self reenactment task. Specifically, NeRFace [16] uses a deep MLP to fit an expression condtioned dynamic NeRF. The current SOTA method HAvatar [69] introduces 3DMM template prior and uses a deep convolutional network to generate a human head NeRF represented by three planes from a mesh template with expression. Note, HAvatar leverages the GAN framework using the adversarial loss function to force the network to generate details that are not view-consistent. For a fair comparison, we remove this part and use VGG perceptual loss as Sec. 4.2 instead.
@@
Next, we conduct a quantitative evaluation for the four methods on 5 identities and 6 cameras using the evaluation split. The evaluation metrics include: Peak Signal-to-Noise Ratio (PSNR), Structure Similarity Index (SSIM), Learned Perceptual Image Patch Similarity (LPIPS) [68] and Fréchet Inception Distance (FID) [22]. As the task mainly focuses on the reconstruction of the head, we use face parsing  [^1] https://github.com/zllrunning/face-parsing.PyTorch to remove the body parts in the image to eliminate their impact in the experiment. As shown in Tab. 1, our method demonstrates a slight improvement in PSNR and SSIM compared with previous methods, and a significant improvement in LPIPS and FID, which means that our method can generate more high-frequency details.
And we also qualitatively compare our method with the above SOTA methods on cross-identity reenactment task. As shown in the results, our method is able to synthesize higher-fidelity images with more accurate expression transfer and richer emotions.
@@ 5.3 Ablation Study
Figure 5: Ablation study on the initialization strategies: FLAME-initialization and our geometry-guided initialization. Our strategy ensures that the hair strands away from the head are well reconstructed.
@@
Table 2: Quantitative evaluation results of the other two ablation baselines and ours.
@@
Figure 6: Ablation study on the deformation Modeling approaches: mesh LBS-based deformation and our fully learned deformation. Our approach can learn complex and exaggerated expressions.
Ablation on Deformation Modeling Approaches. We compare our fully learned deformation field with the previous mesh-based deformation (Mesh-Deform). Specifically, we migrate the method in INSTA [73] for controlling the NeRF deformation to our Gaussians. First we fit a 3DMM mesh template. Then, for each Gaussian point, find the closest face on the mesh, and calculate the deformation gradient to estimate the displacement. Qualitative results are shown in the Fig. 6. For some expressions that cannot be captured well by the 3DMM mesh template, our method can learn accurate deformation, thereby achieving the modeling of complex expressions.
Quantitave results are shown in the Fig. 2. Our method outperforms both the two ablation baselines on PSNR, SSIM and LPIPS metrics.
@@ 6 Discussion and Conclusion
Ethical Considerations. Our method is capable of creating artificial portrait videos, which have the potential to disseminate misinformation, influence public perceptions, and erode confidence in media sources. This can result in considerable negative impacts on society. Hence, it’s critical to consider approaches that can reliably distinguish between authentic and forged content.
Limitation. Although our method has taken a substantial step forward in the quality of image synthesis under complex expressions, one previous problem has still not been solved. For the tongue and teeth inside the mouth or long hair, blurring is sometimes produced in our method due to the lack of tracking methods.
Conclusion. In this paper, we propose Gaussian Head Avatar, a novel representation for head avatar reconstruction, which leverages dynamic 3D Gaussians controlled by a fully learned expression deformation. Experiments demonstrate our method can synthesize ultra high-fidelity images while modeling complex and exaggerated expressions. In addition, we propose a well-designed minute-level initialization strategy to ensure the training convergence of the Gaussian Head Avatar. We believe our Gaussian Head Avatar will become the mainstream direction for head avatar reconstruction under sparse view setups in the future.
Supplementary Material
7 Failure Case
For non-face areas, our method inputs the head pose as the condition to control the deformation, which is not able to model the complex dynamic deformation of long hair, resulting in blurred rendering results as shown in Fig. 7.
On the other hand, the reconstructed head avatar cannot make expressions other than those in the training set. Therefore, when the actor’s expression is too exaggerated, our method will output relatively less exaggerated results as shown in Fig. 8.
Figure 7: Failure case: our method can not reconstruct dynamic long hair.
Figure 8: Failure case: our method produce relatively less exaggerated results.

### v2

@@ Abstract
Creating high-fidelity 3D head avatars has always been a research hotspot, but there remains a great challenge under lightweight sparse view setups. In this paper, we propose Gaussian Head Avatar represented by controllable 3D Gaussians for high-fidelity head avatar modeling. We optimize the neutral 3D Gaussians and a fully learned MLP-based deformation field to capture complex expressions. The two parts benefit each other, thereby our method can model fine-grained dynamic details while ensuring expression accuracy. Furthermore, we devise a well-designed geometry-guided initialization strategy based on implicit SDF and Deep Marching Tetrahedra for the stability and convergence of the training procedure. Experiments show our approach outperforms other state-of-the-art sparse-view methods, achieving ultra high-fidelity rendering quality at 2K resolution even under exaggerated expressions. Project page: https://yuelangx.github.io/gaussianheadavatar.
@@ 1 Introduction
To overcome this bottleneck and further improve the avatar quality, we introduce 3D Gaussian splatting [26] for 3D head avatar modeling. This is an explicit discrete representation that can be well adapted to the rasterization-based rendering pipeline. It has been verified that the 3D Gaussian representation is capable of rendering complex scenes with low computational consumption. Compared to NeRF, the reconstruction quality of static and dynamic scenes [59, 40, 67] is much better while rendering time cost has been significantly reduced. Motivated by this progress, we propose Gaussian Head Avatar, a novel representation that utilizes 3D Gaussian splatting for ultra high-fidelity head avatar modeling. Although recent 4D Gaussian works [59, 67, 40, 66] have been proposed to reconstruct dynamic scenes, all of them cannot be animated. For modeling the animatable head avatar, it is crucial but still unexplored how to effectively control the deformation of 3D Gaussians and model the dynamic appearances through expression coefficients.
@@ 2 Related Works
3D Head Avatar Reconstruction. Due to the wide application value in the film and digital human industry, 3D head avatar reconstruction from multi-view images has always been a research hotspot. Traditional works [31, 4, 19, 8] reconstruct the scan geometry through multi-view stereo and then register a face mesh template to it. However, such methods usually require heavy computation. With the utilization of deep neural networks, current methods [33, 7, 61, 65] achieve very fast reconstruction, producing even more accurate geometry. Lombardi et al. [37], Bi et al. [5] and Ma et al. [41] represent the full head mesh through a deep neural network and train it with multi-view videos as supervision. However, due to the errors in geometric estimation, mesh-based head avatars typically suffer from texture blur. Therefore, some recent methods [57, 38] utilize NeRF representation [43] to synthesize novel view images without geometry reconstruction, or build NeRF on the head mesh template [39]. Furthermore, the NeRF-based methods are extended to sparse view reconstruction tasks [71, 42, 48, 28] and achieve impressive performance.
@@
3D head avatars reconstruction from monocular videos is also a popular yet challenging research topic. Early methods [12, 13, 25, 24, 15, 45] optimize a morphable mesh to fit the training video. Recent methods [20, 27] leverage neural networks to learn non-rigid deformation upon 3DMM face templates [32, 18], thus can recover more dynamic details. Such methods are not flexible enough to handle complex topologies. Therefore, the latest methods explore to construct head avatar models based on implicit SDF [73], point clouds [74] or NeRF [21, 36, 16, 3, 2, 17, 63, 76, 64, 47].
Point-based Rendering. Point elements as a discrete and unstructured representation can fit geometry with arbitrary topology [69] efficiently. Recent methods [58, 30, 29] open up a differentiable rasterization pipeline, such that the point-based representation is widely used in multi-view reconstruction tasks. Aliev et al. [1] and Ruckert et al. [49] propose to first render the feature map, which is transferred to the images through a convolutional renderer. Xu et al.[62] use neural point cloud associated with neural features to model a NeRF. Recently, 3D Gaussian splatting [26] shows its superior performance, beating NeRF in both novel view synthesis quality and rendering speed. Some approaches [59, 67, 40, 66, 34, 72, 72, 50] extend Gaussian representation to dynamic scene reconstruction. However, these methods can not be migrated to the head avatar reconstruction tasks.
@@ 3 Overview
The pipeline of the reconstruction of Gaussian Head Avatar is illustrated in Fig. 2, including the initialization stage and the training stage of Gaussian Head Avatar. Before the beginning of the pipeline, we remove the background [35] of each image and jointly estimate the 3DMM model [18], 3D facial landmarks and the expression coefficients for each frame. In the initialization stage (Sec. 4.3), we reconstruct an SDF-based neutral geometry, and optimize a deformation MLP and a color MLP from the training data as the guidance model. Next, we extract the neutral mesh through DMTet to initialize the neutral Gaussians while the deformation and color MLPs are also inherited from the initialization stage. In the training stage of Gaussian Head Avatar (Sec. 4.2), we deform the neutral Gaussians to the target expression through the dynamic generator given the driving expression coefficients as the condition. Finally, given a camera view, the expressive Gaussians are rendered to a feature map, which is then fed into the convolutional super resolution network to generate high-resolution avatar images. The whole model is optimized under the supervision of multi-view RGB videos.
@@ 4.1 Avatar Representation
$\lambda_{exp}(\cdot)$ and $\lambda_{pose}(\cdot)$ represent the extent to which the points are affected by the expression or the head pose respectively. Without decoupling but just using the expression coefficients as the global condition [16] which also controls the shoulders and upper body, will produce jittering results during the animation. Here, we assume that the Gaussian points closer to 3D landmarks are more affected by the expression coefficients and less affected by the head pose, while the opposite is true for the Gaussian points far away. Specifically, The 3D landmarks $\boldsymbol{P}_{0}$ of the canonical model are first estimated through the 3DMM model in the data preprocessing and then optimized in the initialization stage 4.3. Then for each Gaussian point, we calculate the above weight $\lambda_{exp}(\cdot)$ and $\lambda_{pose}(\cdot)$ as follows:
@@ 4.2 Training
During training, we jointly optimize all the learnable parameters mentioned above in bold, including the neutral Gaussians: $\{\boldsymbol{X}_{0},\boldsymbol{F}_{0},\boldsymbol{Q}_{0},\boldsymbol{S}_{0},\boldsymbol{A}_{0}\}$, the dynamic generator: $\{\boldsymbol{f}_{col}^{exp},\boldsymbol{f}_{col}^{pose},\boldsymbol{f}_{def}^{exp},\boldsymbol{f}_{def}^{pose},\boldsymbol{f}_{att}^{exp},\boldsymbol{f}_{att}^{pose}\}$, and the super resolution network $\boldsymbol{\Psi}$. For the loss function, we only use the foreground RGB images $I_{gt}$ as supervision to construct an L1 loss and a VGG perceptual loss [70] between the generated images $I_{hr}$ and the ground truth $I_{gt}$. Besides, we encourage the first three channels of the 32-channel feature image $I_{C}$ to be RGB channels, which is ensured by a L1 loss term. The total loss is:
@@
with $I_{lr}$ denoting the first three channels of the 32-channel image $I_{C}$. We set the weights $\lambda_{vgg}=0.1$ and $\lambda_{lr}=1$.
@@ 4.3 Geometry-guided Initialization
Unlike neural networks, the Gaussians act as an unordered and unstructured representation. Random initialization leads to failure to converge while naively using the FLAME model to initialize will significantly reduce the reconstruction quality. In this section, we describe in detail how to optimize a mesh guidance model to provide reliable initialization for the Gaussians in Sec. 4.1.
Mesh Guidance Model. Specifically, we first construct a MLP $\boldsymbol{f}_{sdf}$ to represent a signed distance field. In addition, this network will also output the corresponding feature vector of each point, which is used for predicting the point color. It can be formulated as:
@@
Parameters Transfer. Finally, we use the roughly trained mesh guidance model to initialize the Gaussian model. Specifically, we extract the neutral mesh with vertices $X$ and per-vertex features $F$ through DMTet, and directly assign their values to the neutral positions $\boldsymbol{X}_{0}=X$ and the per-vertex feature vectors $\boldsymbol{F}_{0}=F$ of the neutral Gaussians respectively. Then, we retain all the four optimized MLPs: $\{\boldsymbol{f}_{col}^{exp},\boldsymbol{f}_{col}^{pose},\boldsymbol{f}_{def}^{exp},\boldsymbol{f}_{def}^{pose}\}$ for the Gaussian model. For the other neutral attributes: rotation, scale and opacity, we adopt the original initialization strategy in Gaussian Splatting [26]. And the parameters of the two attribute MLPs: $\{\boldsymbol{f}_{att}^{exp},\boldsymbol{f}_{att}^{pose}\}$ and the super resolution network $\boldsymbol{\Psi}$ are just randomly initialized.
@@ 5.1 Implementation Details
Table 1: Quantitative evaluation results of NeRFBlendShape [17], NeRFace [16], HAvatar [71], our method without super resolution and our full method on self reenactment task.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NeRFBlendShape 25.43 0.812 0.148
NeRFace 26.65 0.825 0.151
HAvatar 27.13 0.880 0.65
Ours 27.58 0.882 0.059
Table 2: Quantitative evaluation results of the other SOTA methods and our method on 3D consistency.
Figure 5: Novel view synthesis results of our method. We use 8-view synchronized videos for training the avatar.
@@ 5.2 Results and Comparisons
Self Reenactment. In this section, we first compare our method with existing SOTA methods in qualitative experiments on self reenactment task. Specifically, NeRFace [16] uses a deep MLP to fit an expression condtioned dynamic NeRF. The current SOTA method HAvatar [71] introduces 3DMM template prior and uses a deep convolutional network to generate a human head NeRF represented by three planes from a mesh template with expression. Note, HAvatar leverages the GAN framework using the adversarial loss function to force the network to generate details that are not view-consistent. For a fair comparison, we remove this part and use VGG perceptual loss as Sec. 4.2 instead.
@@
Next, we conduct a quantitative evaluation for the four methods on 5 identities and 6 cameras using the evaluation split. The evaluation metrics include: Peak Signal-to-Noise Ratio (PSNR), Structure Similarity Index (SSIM), Learned Perceptual Image Patch Similarity (LPIPS) [70] and Fréchet Inception Distance (FID) [22]. Note, we calculate FID by comparing the distribution of all the training images and all the rendered images. As the task mainly focuses on the reconstruction of the head, we use face parsing  [^1] https://github.com/zllrunning/face-parsing.PyTorch to remove the body parts in the image to eliminate their impact in the experiment. As shown in Tab. 1, our method demonstrates a slight improvement in PSNR and SSIM compared with previous methods, and a significant improvement in LPIPS and FID, which means that our method can generate more high-frequency details.
Cross-Identity Reenactment. We also qualitatively compare our method with the above SOTA methods on cross-identity reenactment task. As shown in the results, our method is able to synthesize higher-fidelity images with more accurate expression transfer and richer emotions.
Novel View Synthesis. In this section, we show the results of novel view synthesis as shown in Fig. 5 shown. In this case, we use video data from 8 views for training and render the image at a new viewpoint. Next, we quantitatively evaluate the 3D consistency of our method and compare it with other SOTA methods mentioned above. Specifically, we select the 5 identities as above and use the video data from 8 cameras for training while the other 8 hold-out cameras for evaluation. The evaluation metrics include PSNR, SSIM and LPIPS (512 resolution). The results are shown in Tab. 2. Our method outperforms other methods in 3D consistency.
@@ 5.3 Ablation Study
Figure 6: Ablation study on the initialization strategies: FLAME-initialization and our geometry-guided initialization. Our strategy ensures the hair strands away from the head are well reconstructed.
@@
Table 3: Quantitative evaluation results of the other two ablation baselines and ours on self reenactment task.
@@
Figure 7: Ablation study on the deformation modeling: mesh LBS-based deformation and our fully learned deformation. Our approach can learn complex and exaggerated expressions.
Ablation on Deformation Modeling Approaches. We compare our fully learned deformation field with the previous mesh-based deformation (Mesh-Deform). Specifically, we migrate the method in INSTA [76] for controlling the NeRF deformation to our Gaussians. First we fit a 3DMM mesh template. Then, for each Gaussian point, find the closest face on the mesh, and calculate the deformation gradient to estimate the displacement. Qualitative results are shown in the Fig. 7. For some expressions that cannot be captured well by the 3DMM mesh template, our method can learn accurate deformation, thereby achieving the modeling of complex expressions. Quantitave results are shown in the Fig. 3. Our method outperforms both the two ablation baselines on PSNR, SSIM and LPIPS metrics.
@@ 6 Discussion and Conclusion
Ethical Considerations. Our method is capable of creating artificial portrait videos, which have the potential to disseminate misinformation, influence public perceptions, and erode confidence in media sources.
Limitation. For the tongue and teeth inside the mouth or long hair, blurring is sometimes produced in our method due to the lack of tracking methods.
Conclusion. In this paper, we propose Gaussian Head Avatar, a novel representation for head avatar reconstruction, which leverages dynamic 3D Gaussians controlled by a fully learned expression deformation. Experiments demonstrate our method can synthesize ultra high-fidelity images while modeling exaggerated expressions. In addition, we propose a well-designed minute-level initialization strategy to ensure the training convergence. We believe our Gaussian Head Avatar will become the mainstream direction for head avatar reconstruction in the future.
Acknowledgement. This paper is supported by National Key R&D Program of China (2022YFF0902200), the NSFC project No.62125107, the Beijing Municipal Science & Technology Z231100005923030.